OpenStudy (anonymous):

A projectile is thrown upward so that its distance above the ground after t seconds is given by the function h(t) = - 16t2 + 704t. After how many seconds does the projectile take to reach its maximum height? Show your work for bette understandng please!!

OpenStudy (anonymous):

Quadratic equation in standard form \[ax^2 +bx+c\] a=-16 b=704 c=0 Find the maximum value x value->\[\frac{ -b }{ 2a }\] \[\frac{ -704 }{ 2(-16)}\]=22 seconds To find maximum height, plug in x value (time). -16(22^2)+704(22)=7744 ft